OpenStudy (anonymous):

if a spherical tank 2 m in diameter can be filled with a liquid for $300, find the cost to fill the tanks of 6 diameter.

OpenStudy (anonymous):

cost to fill a 6m is 27 * $300 = $8100 The ratio of the volume of a 6m diameter sphere to that of a 2m diameter sphere is 27.\[\frac{36 \pi }{\frac{4 \pi }{3}}=27 \]